Two coils A and B are connected in series across a 110V, 60Hz supply. The resistance of coil A is 2  and inductance of coil B is 0.004H. If the input from the supply is 2.4kW and 1.8kVAR, find the inductance of coil A and resistance of coil B. Also calculate the voltage across each coil.


1
Expert's answer
2021-02-25T11:25:49-0500


The kVA triangle and the circuit is shown above. The circuit kVA is given by, kVA="\sqrt{(2.4)^2+(1.8)^2}=3" or VA = "3000" voltmeters

Circuit current = "\dfrac{3000}{110}=27.27" A

"\therefore27.27^2(R_A+R_B)=2400"

"\Rightarrow R_A+R_B=\dfrac{2400}{742.38}=3.22\Omega"

"\Rightarrow R_B=1.22\Omega"


Impedance of the whole circuit is given by "Z=110/27.27=4.033\Omega"

"X_A+X_B=\sqrt{Z^2-(R_A+R_B)^2}=2.42\Omega"


Now, "X_B=2\pi\times60\times0.004=1.50\Omega"

"\therefore X_A=0.912\Omega"

"\Rightarrow2\pi\times60\times L_A=0.912"

"\Rightarrow L_A=0.00241H"


Now, "Z_A=\sqrt{R_A^2+X_A^2}=2.198\Omega"

P.d across coil A = "I\times Z_A=59.9V"


Now, "Z_B=\sqrt{R_B^2+X_B^2}=1.933\Omega"

P.d across coil B = "I\times Z_B=52.72\Omega"